OpenStudy (anonymous):

Solve log 104x−3 = 7.

OpenStudy (anonymous):

0.096

OpenStudy (anonymous):

Solve log 10^(4x−3) = 7.

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

ok . I got it

OpenStudy (anonymous):

\[\log_{b}a=c --> b^c=a \]

OpenStudy (anonymous):

in your case 10^7=104x-3

OpenStudy (anonymous):

I guess that might be more useful

OpenStudy (anonymous):

thanks for your answer

OpenStudy (anonymous):

because I want take for a test of college

OpenStudy (anonymous):

what test?

OpenStudy (anonymous):

I want study in delta college

OpenStudy (anonymous):

and I must take a test

OpenStudy (anonymous):

placement test?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm VNese

OpenStudy (anonymous):

I just come here for 6month

OpenStudy (anonymous):

VNese?

OpenStudy (anonymous):

vietnamese

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

math is easier here than in asia

OpenStudy (anonymous):

so I need everyone in Open study help me for some question about a math

OpenStudy (anonymous):

that right

OpenStudy (anonymous):

Some people here are very helpful while other just give out answer in hope of getting medal

OpenStudy (anonymous):

what's medal

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

You will see "Good answer" in the right when you click it, you award that person a medal for their answer

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

you're American

OpenStudy (anonymous):

Solve log 10^(4x−3) = 7. =>(4x-3)log10=7 =>4x-3=7, as log10=1 =>x=10/4=2.5